# Mamba Modulation On the Length Generalization of Mamba

### v2

@@ 1 Introduction
In the new age of deep learning, the Transformers [136] architecture has spurred a new age of research into large language models (LLMs) [31, 149, 25, 97, 157] that has largely dominated the space of natural language processing (NLP) research since their introduction. Their surprising capabilities and rapid development have led to their wide application across various domains, including chatbots, intelligent agents, code assistants, etc. However, the Transformer comes with various deficiencies, which has led to research into alternative paradigms that seek to resolve these outstanding concerns. One of these competitors, Mamba [45, 22], is based off the state-space model (SSM) paradigm from control theory [47, 49] that have enabled the training of recurrent models that have overcome the sequential bottleneck of traditional models [116, 55, 15].
@@
Among the various motivation for Mamba and its successors is the goal of length extrapolation, whereby a model that is initially trained on a limited context length (e.g. 2048 tokens in each training sequence) is capable of generalizing to longer sequences at test time (i.e. without further training) due to a more efficient inference-time token processing methodology. However, various works [56, 29, 58] have brought about challenges to this claim. Meanwhile, a key component in the Transformer is the position embedding, for which Rotary Position Embeddings (RoPE) [128] has been a popular choice and applied in many LLMs. Various works have studied RoPE [147, 87] and shown it to be intuitive to manipulate to extend the context window within Transformers [14, 11, 35, 101], whereas no equivalent method yet exists for Mamba-style models. A common explanation for the ability to extend this context length is through as avoiding out-of-distribution (OOD) rotation angles [87, 52] in RoPE, meaning the extended context length (OOD) can be mapped to the in-distribution (ID) context length on which models have been properly trained. However, Mamba does not utilize knowledge of token positions during training, thus making such methods broadly inapplicable.
@@
Recent works [154, 9, 3] have meanwhile made attempts at exploring how to conduct length generalization with Mamba models. A shared feature among these is the focus on a specific input-dependent parameter, $\Delta$, which is used to discretize the underlying state-space model and control for both the state decay over time as well as the incoming input contribution to the state. Their observations rely on the implicit notion that since the duration of the time-step influences the state, it can act as a proxy to filter out (or ignore) parts of the input, or be scaled to influence the long-term information decay within the model state. However, despite the compelling intuition behind such a notion, there remains no fundamental theoretical justification for this.
@@ 2.1 Language Models and Long Contexts
Being capable of modeling long sequences is an important desiderate in various LLM applications. However, due to the quadratic complexity (relative to the sequence length) of the self-attention mechanism in Transformers, long sequence modeling requires a large computational overhead [132]. Early work in efficient Transformers [78, 155, 8, 138, 16] attempted to reduce the computational complexity of attention by inducing greater sparsity. Additional work has explored the use of linear attention [76, 90, 151, 152, 159, 91] to remove the softmax activation that induces this quadratic complexity, however, such methods may impair performance on tasks that demand precise contextual recall [1, 140]. Furthermore, hardware optimizations for more efficient computation [22, 21, 121, 84] as well as inference-time acceleration methods [146] to reduce the computational and memory complexity of Transformers. However, a broader class of linear recurrent models [49, 45, 98, 7, 108, 109], which resemble traditional recurrent neural networks but provide an additional benefit of parallel training over the sequence elements, have emerged as an alternative for long sequences through a sub-quadratic complexity relative to sequence length as well as constant-time inference complexity.
@@ 2.2 Length Generalization and Extrapolation
Yet because many existing methods relevant to Transformer length extrapolation rely explicitly on positional information, it remains an open work to find ways to enable such linear recurrent models to generalize beyond their training lengths. Alternatively, recent works [9, 154, 3] have investigated the post-hoc length extension in Mamba models, with a particular focus on using the discretization time-steps $\Delta_{t}$ for context extension. Ben-Kish et al. [9] use the value of these time-steps to ’decimate’ or remove tokens from the processing of the sequence at specific layers, resulting in a shortened sequence length. Similarly, Ye et al. [154] use the value of the time-steps to filter out tokens. Azizi et al. [3] meanwhile calibrate scaling factors for these time-steps to adjust the long-term decay within the model, extending the context length. Unlike these works, we do not analyze the effect of discretization $(\Delta)$ on extrapolation ability. Instead, we focus on establishing a connection between the spectral characteristics of the state transition matrix and the asymptotic convergence behavior of the hidden state as the input length approaches infinity.
@@ 2.3 Spectrum Analysis of Linear Recurrent Models
Previous works have provided specific analysis of the eigenvalue spectrum of linear recurrent models as a way of understanding their state dynamics and the downstream influence this can have on performance. Gu et al. [46] initially provided an understanding of the specific parameterization of the state transition matrix in SSMs, determining the necessity of a Hurwitz matrix for effective sequence modeling. Orvieto et al. [98] further demonstrated how the eigenvalues have a specific influence on state decay as well as long-term dynamics during training. Beck et al. [7] further bound the state of the recurrence, implicitly bounding the spectrum as well. Finally, Grazzi et al. [44] also recently demonstrate the importance of negative eigenvalues for state-tracking tasks.
@@ 3.1 State-Space Models (SSMs) and Mamba
The SSM-based models, i.e., structured state space sequence models (S4) [49] and Mamba [45] are inspired by the continuous system, which maps a 1-D function or sequence ${\bm{x}}(t)\in\mathbb{R}^{d_{m}}$ to an output ${\bm{y}}(t)\in\mathbb{R}^{d_{m}}$ through a hidden state ${\bm{h}}(t)\in\mathbb{R}^{d_{h}}$. The system uses evolution parameters ${\bm{A}}\in\mathbb{R}^{d_{h}\times d_{h}}$, ${\bm{B}}\in\mathbb{R}^{d_{h}\times d_{m}}$, and ${\bm{C}}\in\mathbb{R}^{d_{m}\times d_{h}}$, creating a continuous system whose dynamics are governed by
@@
The Mamba model uses the selective SSM blocks, which leverage the input-dependent discretization into the recurrence computation. This is done by including an input-dependent timescale parameter $\bm{\Delta}({\bm{x}}_{t})$ to transform the continuous parameters $\bm{A}$}, $\bm{B}$} to discrete parameters $\overline{{\bm{A}}}_{t}$ and $\overline{{\bm{B}}}_{t}$. We follow the official implementation of Mamba [45]:
${\bm{h}}_{t+1}=\overline{\bm{A}}_{t}{\bm{h}}_{t}+\overline{\bm{B}}_{t}{\bm{x}}_{t},\,\quad\,{\bm{y}}_{t}={\bm{C}}_{t}{\bm{h}}_{t},$ (2)
@@
$\overline{{\bm{A}}}_{t}=\exp\left(-\Delta_{t}\odot{\bm{A}}\right),\,\quad\,\overline{{\bm{B}}}_{t}=\Delta_{t}\otimes\left(\left(\Delta{\bm{A}}\right)^{-1}\left(\exp\left(\Delta{\bm{A}}\right)-{\bm{I}}\right)\Delta{\bm{B}}\right),$ (3)
@@
where $\mathsf{Linear}_{\Delta}$, $\mathsf{Linear}_{B}$ and $\mathsf{Linear}_{C}$ are regular linear projections, $\odot$ is the Hadamard product, $\otimes$ is the outer product, $\bm{\Delta}_{t}\in\mathbb{R}^{d}_{+}$ and $\bm{A}=\mathsf{diag}\left(\alpha_{1},\dots,\alpha_{d}\right)\,\text{s.t.}\,\alpha_{i}>0\,\forall i\in\{1,\dots,d\}$. In Mamba, $\Delta$, $\bm{B}$ and $\bm{C}$ are input-dependent, such that at each time-step unique transition matrices can be used to update the system ($\bm{A}$} is left as a fixed parameter in as the dynamics of the state should be consistent across steps). This is based on the observation that some elements in a discrete sequence may not be as important as others, therefore there is an incentive to possibly update the system differently based on this factor. This results in unique update matrices at each time-step $\left(\Delta_{t},\overline{{\bm{A}}}_{t},\overline{{\bm{B}}}_{t},{{\bm{C}}}_{t}\right)$, enabling the ability to solve problems that require selective processing of the sequence. In order to maintain computational efficiency $\bm{A}$} is restricted to having a diagonal structure such that only the diagonal elements of these matrices need to be stored. Mamba2 [22] further restricts the diagonal matrix to have the form of a scalar-times-identity matrix, enabling further computational improvements.
@@ 3.2 Limitations of Mamba in Long Context
Previous work [9, 154, 3] has primarily focused on analyzing the discretization step $\Delta_{t}$ for long context, particularly the vanishing effect of the accumulated term $\exp\left(-\sum_{t=1}^{N}\Delta_{t}\right)$ when $N$ is large and propose different solutions to overcome this out-of-distribution (OOD) issue. For instance,  Azizi et al. [3] propose applying scalar values $s\leq 1$ across different model layers to mitigate OOD discretization steps, ensuring smaller $\Delta_{t}$ to prevent the vanishing issue of distant inputs. They introduce two calibration methods and demonstrate superior length generalization performance in calibrated Mamba models with unconstrained scaling factors. However, their work does not explain why some of resulted scaling factors $s>1$ could still enhance generalization performance.
@@ 4.2 State Convergence in SSMs for Long Contexts
Under the setting of Theorem˜4.2 (proofs in Appendix A), we consider two cases (Mamba and Mamba2) corresponding to the architectural variants of structured state-space models, each characterized by a different form of the structured transition matrix.
@@ 5.1 A Simple Case Analysis on Constant Scaling
To confirm our intuition, we first attempt a simple comparison between the effects of scaling $\Delta_{t}$ and $\bm{A}$}. Here, across all layers, we use a fixed, constant-valued scaling factor. We evaluate language modeling perplexity on the ProofPile dataset [40], following Peng et al. [101], across a varying number of context lengths. This method uses no tuning or training; the scaling is applied explicitly during the forward pass. Figure˜2 shows these results after applying a scaling on perplexity on context lengths from 2K to 128K tokens, with scaling factors of 2 or 3 applied.
@@ 6.1 Language Modeling Perplexity
We first experiment by measuring language modeling perplexity after calibrating scaling factors for either $\Delta_{t}$ or $\bm{A}$}. In this task, we use the black-box zeroth-order calibration method suggested by Azizi et al. [3]; we train a single scaling factor $s_{i}\in\mathbb{R}_{+}$ for every layer $i$ in the model. For a $L$-layer model, this means $L$ individual scaling factors are used. To calibrate, 20 samples of the corresponding context length are used. For example, for a length of $16\mathsf{K}$, 20 samples of this length are used for the calibration of the set of $s_{i}$. Figure˜3 shows these perplexity results on a number of validation datasets, namely ProofPile [40], PG19 [110] and GovReport [63].
@@ 6.2 Passkey Retrieval
Next, we conduct experiments on the Passkey Retrieval task, also known as the Needle-in-A-Haystack. Similar to before, we conduct this to compare the effectiveness of tuning scaling factors for $\bm{A}$} as opposed to $\Delta_{t}$; we again conduct this experiment across different Mamba models. Unlike the language modeling perplexity task however, we train the model on a training set. This training set contains samples of length 4096 corresponding to the task, where the objective is standard instruction-tuning [32]. However, we freeze all parameters except the scaling parameters for each layer. For Mamba, it is equivalent to the number of inner state dimensions, i.e. each inner state utilizes the same scaling factor for each dimension of the SSM state. For Mamba2, this is the number of heads, meaning that each head shares the same scaling factor for each component of its state. Evaluation is conducted on a set of fixed lengths and depths to evaluate for both generalization ability as well as potential biases to relative location within the sequence. The exact setup follows from Ben-Kish et al. [9], in particular, the task comprises of a 5-digit code embedded at a random sequence depth within samples from the WikiText-103 dataset [93]. Models are deemed to have solved length/depth pair if they can correctly solve all evaluation examples, i.e. retrieve the code within the example.
@@ 6.3 LongBench
LongBench [6] is a popular benchmark for testing the long-context abilities of LLMs, serving as a more suitable real-world benchmark on which we can explore how the scaling of $\bm{A}$} as opposed to $\Delta_{t}$ can influence performance. Here, we again use the zeroth-order optimization method as we used for our initial perplexity experiments. More specifically, a constant scaling factor is used for each individual layer. We compare against both the initial base model, as well as MambaExtend. Table˜2 shows results on mamba2-2.7b. In particular, we show that we can increase performance by over 6% through the calibrated scaling of $\bm{A}$}, with a relative improvement of nearly 10% compared to if the scaling was instead calibrated for $\Delta_{t}$.
Table 2: Results on LongBench [6].

### v3

# On the Length Generalization of Mamba

@@ 1 Introduction
In the new age of deep learning, the Transformers [vaswani_attention_2017] architecture has spurred a new age of research into large language models (LLMs) [dubey_llama_2024, yang_qwen25_2024, deepseek-ai_deepseek-v3_2024, openai_gpt-4_2023, zeng_zeta_2025] that has largely dominated the space of natural language processing (NLP) research since their introduction. Their surprising capabilities and rapid development have led to their wide application across various domains, including chatbots, intelligent agents, code assistants, etc. However, the Transformer comes with various deficiencies, which has led to research into alternative paradigms that seek to resolve these outstanding concerns. One of these competitors, Mamba [gu_mamba_2024, dao_transformers_2024], is based off the state-space model (SSM) paradigm from control theory [gu_combining_2021, gu_efficiently_2022] that have enabled the training of recurrent models that have overcome the sequential bottleneck of traditional models [rumelhart_parallel_1989, hochreiter_long_1997, cho_learning_2014].
@@
Among the various motivation for Mamba and its successors is the goal of length extrapolation, whereby a model that is initially trained on a limited context length (e.g. 2048 tokens in each training sequence) is capable of generalizing to longer sequences at test time (i.e. without further training) due to a more efficient inference-time token processing methodology. However, various works [hsieh_ruler_2024, dong_bamboo_2024, huang_how_2025] have brought about challenges to this claim. Meanwhile, a key component in the Transformer is the position embedding, for which Rotary Position Embeddings (RoPE) [su_roformer_2024] has been a popular choice and applied in many LLMs. Various works have studied RoPE [xu_base_2024, liu_scaling_2024] and shown it to be intuitive to manipulate to extend the context window within Transformers [chen_extending_2023, bloc97_ntk-aware_2023, emozilla_dynamically_2023, peng_yarn_2024], whereas no equivalent method yet exists for Mamba-style models. A common explanation for the ability to extend this context length is through as avoiding out-of-distribution (OOD) rotation angles [liu_scaling_2024, han_lm-infinite_2024] in RoPE, meaning the extended context length (OOD) can be mapped to the in-distribution (ID) context length on which models have been properly trained. However, Mamba does not utilize knowledge of token positions during training, thus making such methods broadly inapplicable.
@@
Recent works [ye_longmamba_2025, ben-kish_decimamba_2025, azizi_mambaextend_2025] have meanwhile made attempts at exploring how to conduct length generalization with Mamba models. A shared feature among these is the focus on a specific input-dependent parameter, $\Delta$, which is used to discretize the underlying state-space model and control for both the state decay over time as well as the incoming input contribution to the state. Their observations rely on the implicit notion that since the duration of the time-step influences the state, it can act as a proxy to filter out (or ignore) parts of the input, or be scaled to influence the long-term information decay within the model state. However, despite the compelling intuition behind such a notion, there remains no fundamental theoretical justification for this.
@@ 2.1 Language Models and Long Contexts
Being capable of modeling long sequences is an important desiderate in various LLM applications. However, due to the quadratic complexity (relative to the sequence length) of the self-attention mechanism in Transformers, long sequence modeling requires a large computational overhead [tay_long_2021]. Early work in efficient Transformers [kitaev_reformer_2020, zaheer_big_2020, beltagy_longformer_2020, wang_linformer_2020, choromanski_rethinking_2021] attempted to reduce the computational complexity of attention by inducing greater sparsity. Additional work has explored the use of linear attention [katharopoulos_transformers_2020, ssm-pooler, yang_gated_2024, yang_parallelizing_2024, zhang_gated_2024, regla] to remove the softmax activation that induces this quadratic complexity, however, such methods may impair performance on tasks that demand precise contextual recall [arora_zoology_2024, resona]. Furthermore, hardware optimizations for more efficient computation [dao_transformers_2024, dao_flashattention-2_2024, shah_flashattention-3_2024, liu_ringattention_2024] as well as inference-time acceleration methods [xiao_efficient_2024] to reduce the computational and memory complexity of Transformers. However, a broader class of linear recurrent models [gu_efficiently_2022, gu_mamba_2024, orvieto_resurrecting_2023, beck_xlstm_2024, qin_hierarchically_2023, qin_hgrn2_2024], which resemble traditional recurrent neural networks but provide an additional benefit of parallel training over the sequence elements, have emerged as an alternative for long sequences through a sub-quadratic complexity relative to sequence length as well as constant-time inference complexity.
@@ 2.2 Length Generalization and Extrapolation
Yet because many existing methods relevant to Transformer length extrapolation rely explicitly on positional information, it remains an open work to find ways to enable such linear recurrent models to generalize beyond their training lengths. Alternatively, recent works [ben-kish_decimamba_2025, ye_longmamba_2025, azizi_mambaextend_2025] have investigated the post-hoc length extension in Mamba models, with a particular focus on using the discretization time-steps $\Delta_{t}$ for context extension. ben-kish_decimamba_2025 use the value of these time-steps to ’decimate’ or remove tokens from the processing of the sequence at specific layers, resulting in a shortened sequence length. Similarly, ye_longmamba_2025 use the value of the time-steps to filter out tokens. azizi_mambaextend_2025 meanwhile calibrate scaling factors for these time-steps to adjust the long-term decay within the model, extending the context length. Unlike these works, we do not analyze the effect of discretization $(\Delta)$ on extrapolation ability. Instead, we focus on establishing a connection between the spectral characteristics of the state transition matrix and the asymptotic convergence behavior of the hidden state as the input length approaches infinity.
@@ 2.3 Spectrum Analysis of Linear Recurrent Models
Previous works have provided specific analysis of the eigenvalue spectrum of linear recurrent models as a way of understanding their state dynamics and the downstream influence this can have on performance. gu_hippo_2020 initially provided an understanding of the specific parameterization of the state transition matrix in SSMs, determining the necessity of a Hurwitz matrix for effective sequence modeling. orvieto_resurrecting_2023 further demonstrated how the eigenvalues have a specific influence on state decay as well as long-term dynamics during training. beck_xlstm_2024 further bound the state of the recurrence, implicitly bounding the spectrum as well. Finally, grazzi_unlocking_2025 also recently demonstrate the importance of negative eigenvalues for state-tracking tasks.
@@ 3.1 State-Space Models (SSMs) and Mamba
The SSM-based models, i.e., structured state space sequence models (S4) [gu_efficiently_2022] and Mamba [gu_mamba_2024] are inspired by the continuous system, which maps a 1-D function or sequence ${\bm{x}}(t)\in\mathbb{R}^{d_{m}}$ to an output ${\bm{y}}(t)\in\mathbb{R}^{d_{m}}$ through a hidden state ${\bm{h}}(t)\in\mathbb{R}^{d_{h}}$. The system uses evolution parameters ${\bm{A}}\in\mathbb{R}^{d_{h}\times d_{h}}$, ${\bm{B}}\in\mathbb{R}^{d_{h}\times d_{m}}$, and ${\bm{C}}\in\mathbb{R}^{d_{m}\times d_{h}}$, creating a continuous system whose dynamics are governed by
@@
The Mamba model uses the selective SSM blocks, which leverage the input-dependent discretization into the recurrence computation. This is done by including an input-dependent timescale parameter $\bm{\Delta}({\bm{x}}_{t})$ to transform the continuous parameters $\bm{A}$}, $\bm{B}$} to discrete parameters $\overline{{\bm{A}}}_{t}$ and $\overline{{\bm{B}}}_{t}$. We follow the official implementation of Mamba [gu_mamba_2024]:
${\bm{h}}_{t}=\overline{\bm{A}}_{t}{\bm{h}}_{t-1}+\overline{\bm{B}}_{t}{\bm{x}}_{t},\,\quad\,{\bm{y}}_{t}={\bm{C}}_{t}{\bm{h}}_{t},$ (2)
@@
$\overline{{\bm{A}}}_{t}=\exp\left(-\Delta_{t}\odot{\bm{A}}\right),\,\quad\,\overline{{\bm{B}}}_{t}=\Delta_{t}\otimes\left(\left(\Delta_{t}{\bm{A}}\right)^{-1}\left(\exp\left(\Delta_{t}{\bm{A}}\right)-{\bm{I}}\right){\bm{B}}\right),$ (3)
@@
where $\mathsf{Linear}_{\Delta}$, $\mathsf{Linear}_{B}$ and $\mathsf{Linear}_{C}$ are regular linear projections, $\odot$ is the Hadamard product, $\otimes$ is the outer product, $\bm{\Delta}_{t}\in\mathbb{R}^{d}_{+}$ and $\bm{A}=\mathsf{diag}\left(\alpha_{1},\dots,\alpha_{d}\right)\,\text{s.t.}\,\alpha_{i}>0\,\forall i\in\{1,\dots,d\}$. In Mamba, $\Delta$, $\bm{B}$ and $\bm{C}$ are input-dependent, such that at each time-step unique transition matrices can be used to update the system ($\bm{A}$} is left as a fixed parameter in as the dynamics of the state should be consistent across steps). This is based on the observation that some elements in a discrete sequence may not be as important as others, therefore there is an incentive to possibly update the system differently based on this factor. This results in unique update matrices at each time-step $\left(\Delta_{t},\overline{{\bm{A}}}_{t},\overline{{\bm{B}}}_{t},{{\bm{C}}}_{t}\right)$, enabling the ability to solve problems that require selective processing of the sequence. In order to maintain computational efficiency $\bm{A}$} is restricted to having a diagonal structure such that only the diagonal elements of these matrices need to be stored. Mamba2 [dao_transformers_2024] further restricts the diagonal matrix to have the form of a scalar-times-identity matrix, enabling further computational improvements.
@@ 3.2 Limitations of Mamba in Long Context
Previous work [ben-kish_decimamba_2025, ye_longmamba_2025, azizi_mambaextend_2025] has primarily focused on analyzing the discretization step $\Delta_{t}$ for long context, particularly the vanishing effect of the accumulated term $\exp\left(-\sum_{t=1}^{N}\Delta_{t}\right)$ when $N$ is large and propose different solutions to overcome this out-of-distribution (OOD) issue. For instance,  azizi_mambaextend_2025 propose applying scalar values $s\leq 1$ across different model layers to mitigate OOD discretization steps, ensuring smaller $\Delta_{t}$ to prevent the vanishing issue of distant inputs. They introduce two calibration methods and demonstrate superior length generalization performance in calibrated Mamba models with unconstrained scaling factors. However, their work does not explain why some of resulted scaling factors $s>1$ could still enhance generalization performance.
@@ 4.2 State Convergence in SSMs for Long Contexts
Under the setting of Theorem˜4.2 (proofs in Appendix˜A), we consider two cases (Mamba and Mamba2) corresponding to the architectural variants of structured state-space models, each characterized by a different form of the structured transition matrix.
@@ 5.1 A Simple Case Analysis on Constant Scaling
To confirm our intuition, we first attempt a simple comparison between the effects of scaling $\Delta_{t}$ and $\bm{A}$}. Here, across all layers, we use a fixed, constant-valued scaling factor. We evaluate language modeling perplexity on the ProofPile dataset [gao_pile_2021], following peng_yarn_2024, across a varying number of context lengths. This method uses no tuning or training; the scaling is applied explicitly during the forward pass. Figure˜2 shows these results after applying a scaling on perplexity on context lengths from 2K to 128K tokens, with scaling factors of 2 or 3 applied.
@@ 6.1 Language Modeling Perplexity
We first experiment by measuring language modeling perplexity after calibrating scaling factors for either $\Delta_{t}$ or $\bm{A}$}. In this task, we use the black-box zeroth-order calibration method suggested by azizi_mambaextend_2025; we train a single scaling factor $s_{i}\in\mathbb{R}_{+}$ for every layer $i$ in the model. For a $L$-layer model, this means $L$ individual scaling factors are used. To calibrate, 20 samples of the corresponding context length are used. For example, for a length of $16\mathsf{K}$, 20 samples of this length are used for the calibration of the set of $s_{i}$. Figure˜3 shows these perplexity results on a number of validation datasets, namely ProofPile [gao_pile_2021], PG19 [rae_compressive_2020] and GovReport [huang_efficient_2021].
@@ 6.2 Passkey Retrieval
Next, we conduct experiments on the Passkey Retrieval task, also known as the Needle-in-A-Haystack. Similar to before, we conduct this to compare the effectiveness of tuning scaling factors for $\bm{A}$} as opposed to $\Delta_{t}$; we again conduct this experiment across different Mamba models. Unlike the language modeling perplexity task however, we train the model on a training set. This training set contains samples of length 4096 corresponding to the task, where the objective is standard instruction-tuning [dubois_alpacafarm_2023]. However, we freeze all parameters except the scaling parameters for each layer. For Mamba, it is equivalent to the number of inner state dimensions, i.e. each inner state utilizes the same scaling factor for each dimension of the SSM state. For Mamba2, this is the number of heads, meaning that each head shares the same scaling factor for each component of its state. Evaluation is conducted on a set of fixed lengths and depths to evaluate for both generalization ability as well as potential biases to relative location within the sequence. The exact setup follows from ben-kish_decimamba_2025, in particular, the task comprises of a 5-digit code embedded at a random sequence depth within samples from the WikiText-103 dataset [merity_pointer_2017]. Models are deemed to have solved length/depth pair if they can correctly solve all evaluation examples, i.e. retrieve the code within the example.
@@ 6.3 LongBench
LongBench [bai_longbench_2024] is a popular benchmark for testing the long-context abilities of LLMs, serving as a more suitable real-world benchmark on which we can explore how the scaling of $\bm{A}$} as opposed to $\Delta_{t}$ can influence performance. Here, we again use the zeroth-order optimization method as we used for our initial perplexity experiments. More specifically, a constant scaling factor is used for each individual layer. We compare against both the initial base model, as well as MambaExtend. Table˜2 shows results on mamba2-2.7b. In particular, we show that we can increase performance by over 6% through the calibrated scaling of $\bm{A}$}, with a relative improvement of nearly 10% compared to if the scaling was instead calibrated for $\Delta_{t}$.
Table 2: Results on LongBench [bai_longbench_2024].